OpenStudy (anonymous):

Which of the following are the x and y intercepts of the equation 4x - 2y = 8? A. (3, 0) and (0, 4) B. (2, 0) and (0, 4) C. (-2, 0) and (0, -4) D. (2, 0) and (0, -4)

OpenStudy (anonymous):

C or D... From there it's tough, dont have any sort of graph.

OpenStudy (anonymous):

ok. when you need to find the y intercept, set the equation equal to y: y=mx+b the b will be your y intercept

OpenStudy (sasogeek):

D it is D because at the X intercept, y = 0 and at the Y intercept, x = 0 when x = 0, -2y=8 y=8/-2 y=-4 =>Y intercept (0,-4) when y = 0 4x = 8 x = 8/4 x = 2 => X intercept (2,0) hence D, (2,0) and (0,-4)